# Zijie Lin1, Yang Zhang1†, Xiaoyan Zhao2, Fengbin Zhu1, Fuli Feng3, Tat-Seng Chua1

### v1

@@ Abstract
Large Language Models (LLMs) have shown strong potential for recommendation by framing item prediction as a token-by-token language generation task. However, existing methods treat all item tokens equally, simply pursuing likelihood maximization during both optimization and decoding. This overlooks crucial token-level differences in decisiveness—many tokens contribute little to item discrimination yet can dominate optimization or decoding. To quantify token decisiveness, we propose a novel perspective that models item generation as a decision process, measuring token decisiveness by the Information Gain (IG) each token provides in reducing uncertainty about the generated item. Our empirical analysis reveals that most tokens have low IG but often correspond to high logits, disproportionately influencing training loss and decoding, which may impair model performance. Building on these insights, we introduce an Information Gain-based Decisiveness-aware Token handling (IGD) strategy that integrates token decisiveness into both tuning and decoding. Specifically, IGD downweights low-IG tokens during tuning and rebalances decoding to emphasize tokens with high IG. In this way, IGD moves beyond pure likelihood maximization, effectively prioritizing high-decisiveness tokens. Extensive experiments on four benchmark datasets with two LLM backbones demonstrate that IGD consistently improves recommendation accuracy, achieving significant gains on widely used ranking metrics compared to strong baselines. Our codes are available at https://github.com/ZJLin2oo1/IGD.
@@ 1 Introduction
Recommendation systems [1, 2] play a crucial role in helping users discover relevant and personalized content. With recent advances in Large Language Models (LLMs) [3, 4, 5], there is growing interest in leveraging LLMs’ strong language understanding and reasoning capabilities [6, 7] for recommendation tasks [8, 9, 10, 11, 12, 13, 14, 15, 16], giving rise to a new paradigm known as LLM4Rec [17, 18]. In this paradigm, recommendation is typically formulated as a natural language problem: user history and task context are encoded into a prompt, based on which the LLM is tuned to generate the top-$K$ recommended items via autoregressive token decoding [19]. This approach has demonstrated strong performance in capturing user intent and generating personalized outputs [20, 10, 21].
Despite its promise, we argue that the existing token handling strategy in LLM4Rec does not fully align with the item generation process in recommendation. The current approach is primarily likelihood-driven, treating each item token equally and simply focusing on: 1) optimizing token likelihood during fine-tuning for data fitting, and 2) maximizing token likelihood during inference for generation [19, 20]. However, not all tokens are equally important. Some tokens are more decisive in defining the item, while others serve grammatical or filler functions with low decisiveness. Low-decisiveness tokens do not reduce uncertainty in item generation, making their focus less meaningful. Moreover, low-decisiveness tokens may have high logits—such as "ghost tokens" defined in [19], which have generation probabilities close to 1 for a given prefix. These tokens can dominate likelihood-maximizing decoding, introducing bias toward items with more such tokens [19]. To improve both tuning and decoding, it is crucial to quantify and incorporate token decisiveness, rather than relying solely on likelihood optimization.
@@ 3 Token Decisiveness Modeling
[…]
@@ 3.1 Token Decisiveness Measurement
$H(y_{\leq t})=-\sum_{\mathcal{I}_{i}\in\mathcal{I}^{y_{\leq t}}}p_{r}(\mathcal{I}_{i})\log p_{r}(\mathcal{I}_{i})$ (4)
@@ 5.1 Experimental Setup
For standard recommendation settings, we compare our method against: (1) two representative sequential recommendation models—GRU4Rec[27] and SASRec[28]; and (2) two state-of-the-art (SOTA) LLM-based recommendation approaches—BIGRec[20] and D3[19]. Our IGD strategy can be integrated into both BIGRec and D3 for fair comparison. To further evaluate the effectiveness of IGD as a token handling strategy, we compare it with two token handing baselines that can also be seamlessly integrated into LLM4Rec frameworks: 1) Position Normalization (Pos), which assigns higher weight to earlier item tokens to mitigate position bias; and 2) Causal Fine-tuning (CFT) [29], which builds upon Pos by introducing an additional causal loss term to enhance the modeling of causal effects at the token level. See Appendix A for detailed descriptions of all baselines. By default, we implement all LLM-based methods using Qwen2.5-1.5B [25]. More implementation details, including hyperparameter tuning settings, can be found in Appendix B.
@@ 5.2 Main Results (RQ1)
In this section, we evaluate whether the proposed IGD method improves overall recommendation performance. Table 2 summarizes the performance of all compared methods. For token handling strategies (IGD, CFT, and Pos), we implement each on top of both BIGRec and D3 for comparison. For traditional baselines such as GRU4Rec and SASRec, we adopt the reported results from the D3 paper [19] to ensure consistency. Our key observations are as follows:
@@
  • LLM4Rec methods (BIGRec, D3) clearly outperform traditional recommendation models (GRU4Rec, SASRec). This performance gap becomes even larger when combined with token reweighting techniques, highlighting the advantages of leveraging LLMs in recommendation tasks.
@@ 5.5 Generalizability (RQ4)
To demonstrate the generalizability of our proposed IGD method, we evaluate it under different tokenization strategies and model scales. Specifically, we conduct experiments using the LLaMA3-8B model [32], with the prefix tree constructed using the LLaMA3 tokenizer [32]. To enable efficient fine-tuning, we employ 4-bit QLoRA [33] for both training and inference, with the rank parameter $r$ set to 32, the scaling factor $\alpha$ set to 64, and the dropout rate set to 0.05. As shown in Table 4, IGD yields consistent improvements across all four datasets when applied to D3, with an average gain of 26.8% in HR@10 and 26.4% in NDCG@10. These results confirm the generalizability of our IGD approach across different model architectures.
@@ 6 Related Work
Token-level Biases in LLM4Rec: Autoregressive LLMs generate items token-by-token, misaligning with item-level recommendation objectives [42, 43, 19]. Multi-token items suffer from token-level biases: high-probability ghost tokens dominate decoding without aiding discrimination, inflating scores for longer items [19]; and common tokens share across many items (e.g., ‘The’) dilute the impact of rare and informative tokens, reducing diversity [42, 19].
@@ 7 Conclusion & Limitation
Our current experiments focus exclusively on IG-based token scoring. Future work may explore alternative decision metrics, such as Gini impurity, Gain Ratio, or Chi-squared statistics [44, 45, 46], which may provide complementary perspectives on token decisiveness.
@@ Appendix B Implementation Details of Compared Method
For traditional baselines, we directly adopt the settings from the D3 paper [19], as our experimental setup is fully aligned with theirs. For LLM-based methods, we use Qwen2.5-1.5B [25] as the backbone. The batch size is set to 64, the optimizer is AdamW, the learning rate is $1\times 10^{-4}$, and the dropout rate is 0.05. Model selection is based on validation loss, with an early stopping strategy that uses a maximum of 3 epochs and a patience of 1 epoch. All other settings follow the D3 paper. For our proposed IGD method, the tuning hyperparameter $\beta$ is selected from the set $\{0.1,0.2,0.3,0.4,0.5,0.6,1.0\}$, and the decoding hyperparameter $\alpha$ is selected from $\{0.0,0.1,0.2,0.3,0.4\}$. We first search for the optimal $\beta$ based on validation performance, followed by a search for the best $\alpha$.
@@ Appendix E All Results for Generalizability Study
This section presents all the results for the Generalizability study. The results are summarized in Table 5

### v3

# Zijie Lin1, Yang Zhang1†, Xiaoyan Zhao2, Fengbin Zhu1†, Fuli Feng3, Tat-Seng Chua1

@@ Abstract
Large Language Models (LLMs) have shown strong potential for recommendation by framing item prediction as a token-by-token language generation task. However, existing methods treat all item tokens equally, simply pursuing likelihood maximization during both optimization and decoding. This overlooks crucial token-level differences in decisiveness—many tokens contribute little to item discrimination yet can dominate optimization or decoding. To quantify token decisiveness, we propose a novel perspective that models item generation as a decision process, measuring token decisiveness by the Information Gain (IG) each token provides in reducing uncertainty about the generated item. Our empirical analysis reveals that most tokens have low IG but often correspond to high logits, disproportionately influencing training loss and decoding, which may impair model performance. Building on these insights, we introduce an Information Gain-based Decisiveness-aware Token handling (IGD) strategy that integrates token decisiveness into both tuning and decoding. Specifically, IGD downweights low-IG tokens during tuning and rebalances decoding to emphasize tokens with high IG. In this way, IGD moves beyond pure likelihood maximization, effectively prioritizing high-decisiveness tokens. Extensive experiments on four benchmark datasets with two LLM bbones demonstrate that IGD consistently improves recommendation accuracy, achieving significant gains on widely used ranking metrics compared to strong baselines. Our codes are available at https://github.com/ZJLin2oo1/IGD.
@@ 1 Introduction
Recommendation systems [1, 2, 3] play a crucial role in helping users discover relevant and personalized content. With recent advances in Large Language Models (LLMs) [4, 5, 6], there is growing interest in leveraging LLMs’ strong language understanding and reasoning capabilities [7, 8, 9] for recommendation tasks [10, 11, 12, 13, 14, 15], giving rise to a new paradigm known as LLM4Rec [16, 17, 18]. In this paradigm, recommendation is typically formulated as a natural language problem: user history and task context are encoded into a prompt, based on which the LLM is tuned to generate the top-$K$ recommended items via autoregressive token decoding [19]. This approach has demonstrated strong performance in capturing user intent and generating personalized outputs [20, 21, 22].
Despite its promise, we argue that the existing token handling strategy in LLM4Rec does not fully align with the item generation process in recommendation. The current approach is primarily likelihood-driven, treating each item token equally and simply focusing on: 1) optimizing token likelihood during fine-tuning for data fitting [23, 24, 21], and 2) maximizing token likelihood during inference for generation [19, 20]. However, not all tokens are equally important. Some tokens are more decisive in defining the item, while others serve grammatical or filler functions with low decisiveness. Low-decisiveness tokens do not reduce uncertainty in item generation, making their focus less meaningful. Moreover, low-decisiveness tokens may have high logits—such as "ghost tokens" defined in [19], which have generation probabilities close to 1 for a given prefix. These tokens can dominate likelihood-maximizing decoding, introducing bias toward items with more such tokens [19]. To improve both tuning and decoding, it is crucial to quantify and incorporate token decisiveness, rather than relying solely on likelihood optimization.
@@ 3 Token Decisiveness Modeling
Figure 1: Illustration of LLM4Rec autoregressive token generation as a sequential decision process. As tokens are generated, the entropy of the remaining sequence gradually decreases. The information gain (IG) quantifies this reduction, e.g., $\text{IG}(\text{M};\text{S})$ measures the IG of token "Mario" given prefix "Super". Tokens shared across many items (e.g., "The") exhibit lower decisiveness with lower IG, while more decisive tokens (e.g., "Super") lead to larger IG. Tokens with IG=0—such as "of", "Zelda", "Sword", and "Man"—are referred to as zero-IG tokens.
@@ 3.1 Token Decisiveness Measurement
$H(y_{\leq t})=-\sum_{\mathcal{I}_{i}\in\mathcal{I}^{y_{\leq t}}}p_{r}(\mathcal{I}_{i})\log p_{r}(\mathcal{I}_{i}),$ (4)
@@ 5.1 Experimental Setup
For standard recommendation settings, we compare our method against: (1) two representative sequential recommendation models—GRU4Rec[30], SASRec[31] and LRURec[32]; and (2) two state-of-the-art (SOTA) LLM-based recommendation approaches—BIGRec[20] and D3[19]. Our IGD strategy can be integrated into both BIGRec and D3 for fair comparison. To further evaluate the effectiveness of IGD as a token handling strategy, we compare it with two token handing baselines that can also be seamlessly integrated into LLM4Rec frameworks: 1) Position Normalization (Pos), which assigns higher weight to earlier item tokens to mitigate position bias; and 2) Causal Fine-tuning (CFT) [33], which builds upon Pos by introducing an additional causal loss term to enhance the modeling of causal effects at the token level. See Appendix A for detailed descriptions of all baselines. By default, we implement all LLM-based methods using Qwen2.5-1.5B [28]. More implementation details, including hyperparameter tuning settings, can be found in Appendix B.
@@ 5.2 Main Results (RQ1)
In this section, we evaluate whether the proposed IGD method improves overall recommendation performance. Table 2 summarizes the performance of all compared methods. For token handling strategies (IGD, CFT, and Pos), we implement each on top of both BIGRec and D3 for comparison. For traditional baselines such as GRU4Rec and SASRec, we adopt the reported results from the D3 paper [19] to ensure consistency. For LRURec, we utilize the official PyTorch implementation [32] and evaluate it using our dataset split. Our key observations are as follows:
@@
  • LLM4Rec methods (BIGRec, D3) clearly outperform traditional recommendation models (GRU4Rec, SASRec). Although LRURec outperforms them on the CDs and Books datasets, BIGRec and D3 regain the lead when combined with token reweighting techniques, highlighting the advantages of leveraging LLMs in recommendation tasks.
@@ 5.5 Generalizability (RQ4)
To demonstrate the generalizability of our proposed IGD method, we evaluate it under different tokenization strategies and model scales. Specifically, we conduct experiments using the LLaMA3-8B model [36], with the prefix tree constructed using the LLaMA3 tokenizer [36]. To enable efficient fine-tuning, we employ 4-bit QLoRA [37] for both training and inference, with the rank parameter $r$ set to 32, the scaling factor $\alpha$ set to 64, and the dropout rate set to 0.05. As shown in Table 4, IGD yields consistent improvements across all four datasets when applied to D3, with an average gain of 26.8% in HR@10 and 26.4% in NDCG@10. In addition to the Amazon dataset, we also evaluate IGD on the Steam dataset (see Appendix E), where it demonstrates consistent improvements. These results confirm the generalizability of our IGD approach across different model architectures and datasets.
@@ 6 Related Work
Token-level Biases in LLM4Rec: Autoregressive LLMs generate items token-by-token, misaligning with item-level recommendation objectives [46, 19, 47, 48]. Multi-token items suffer from token-level biases: high-probability ghost tokens dominate decoding without aiding discrimination, inflating scores for longer items [19]; and common tokens share across many items (e.g., ‘The’) dilute the impact of rare and informative tokens, reducing diversity [46, 19].
@@
Token Prefix Trie for LLM4Rec: Recent works leverage token prefix trie for LLM-based recommendation, each employing the trie in a different way to guide token generation and learning. MSL [48] uses the trie as a constraint: a masked softmax prunes infeasible next tokens, reducing negative optimization signals and focusing training on valid continuations. Flower [47] uses the trie for process-guided supervision: rewards are stored at nodes and propagated along paths during tuning, encouraging trajectories that obtain higher rewards. Ours uses the trie for decisiveness-aware weighting: each node stores an IG score estimated from data, and these scores are applied to debias both tuning and decoding, highlighting discriminative tokens while down-weighting ambiguous ones.
@@ 7 Conclusion & Limitation
Our current experiments focus exclusively on IG-based token scoring. However, the IG values of tokens are highly skewed (see Figure 4), which makes it challenging to design practical reweighting methods (as discussed in Appendix G). Future work may explore alternative decision metrics, such as Gini impurity, Gain Ratio, or Chi-squared statistics [49, 50, 51], which may provide complementary perspectives on token decisiveness.
In addition, our present analysis is restricted to text tokens. Extending decisiveness modeling to semantic ID tokens [34, 52, 53] and multimodal tokens [54, 55, 56] is a promising direction for future research.
8 Acknowledgement
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-002) and A*STAR under its Japan-Singapore Joint Call: Japan Science and Technology Agency (JST) and Agency for Science, Technology and Research (A*STAR) 2024 (Award R24I6IR142). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the A*STAR.
@@ Appendix A The Details of Compared Methods
  • LRURec [32]: A linear recurrent unit-based sequential recommender that enables fast, incremental inference with reduced model size and parallelizable training, achieving strong accuracy and efficiency compared to attention-based baselines.
@@ Appendix B Implementation Details of Compared Method
For traditional baselines, we directly adopt the settings from the D3 paper [19], as our experimental setup is fully aligned with theirs. For LLM-based methods, we use Qwen2.5-1.5B [28] as the backbone. The batch size is set to 64, the optimizer is AdamW, the learning rate is $1\times 10^{-4}$, and the dropout rate is 0.05. Model selection is based on validation loss, with an early stopping strategy that uses a maximum of 3 epochs and a patience of 1 epoch. All other settings follow the D3 paper. For our proposed IGD method, the tuning hyperparameter $\beta$ is selected from the set $\{0.08,0.1,0.2,0.3,0.4,0.5,0.6,1.0\}$, and the decoding hyperparameter $\alpha$ is selected from $\{0.0,0.1,0.2,0.3,0.4\}$. We first search for the optimal $\beta$ based on validation performance, followed by a search for the best $\alpha$.
@@ Appendix E All Results for Generalizability Study
This section presents all the results for the Generalizability study. The results are summarized in Table 5 and Table 6
@@
More dataset.
To further assess generalizability, we additionally evaluated the best-performing baseline (D3) and its IGD-enhanced variant on Steam dataset ($\sim$982K interactions). Results in Table 6 show that IGD continues to provide consistent gains over the baseline.
Table 6: Results on the Steam dataset ($\sim$982K interactions) under the same evaluation protocol as the main study.
Method HR@10 NDCG@10
D3 0.1126 0.0782
+IGD 0.1184 (+5.15%) 0.0810 (+3.58%)
Appendix F Item Diversity under IGD-Decoding
We evaluate IGD-D’s effect on item diversity with $\alpha\in\{0.0,0.1,0.2,0.3,0.4\}$.
First Word Repetition Rate (FWR)—lower is better—is the proportion of the most frequent first token among the top-10 recommended items. Item Score Entropy (ISE)—higher is better—is computed from the decoding probabilities induced by final item scores using the standard $\sum-p\log p$ aggregation.
As Table 7 is shown, increasing $\alpha$ consistently decreases FWR and increases ISE across all datasets.
[…]
CDs Toys Games Books
[…]
0.0 0.364 2.65 0.588 2.67 0.340 2.95 0.382 2.75
0.1 0.332 2.70 0.577 2.73 0.328 2.98 0.339 2.81
0.2 0.304 2.76 0.564 2.80 0.321 3.01 0.306 2.87
0.3 0.282 2.84 0.557 2.88 0.307 3.04 0.278 2.93
0.4 0.265 2.92 0.551 2.95 0.296 3.08 0.259 2.99
Mechanism Analysis: Non-decisive tokens tend to receive high logits and dominate beam expansion, reducing diversity. IGD-D increases the logits of decisive tokens (via $\alpha$), yielding lower FWR and higher ISE.
Appendix G Why IGD-Tuning Adopts a Binary Weighting Scheme
Our binary-based weighting scheme is motivated by the observed behavior of IG in our setting and is not arbitrary. Two empirical observations guide the design:
(1) Distinct zero-IG token group. Tokens with zero IG form a dominant cluster (about 55% of all tokens) and exhibit disproportionately high logits with very low training loss. Treating this group as a separate class and applying a uniform down-weighting factor $\beta$ effectively mitigates their outsized influence.
(2) Non-linear IG distribution among non-zero tokens. The IG distribution for non-zero tokens is highly skewed and roughly exponential rather than linear. This makes it difficult to craft a smooth, well-calibrated continuous weighting function over IG. We experimented with a simple linear and monotonic alternative:
$w_{t}\;=\;\beta\;+\;(1-\beta)\cdot\frac{\mathrm{IG}}{\mathrm{IG}_{\max}}$
As shown in Table 8, this linear function does not outperform the binary scheme.
Table 8: Comparison of linear vs. binary IGD-T weighting across datasets. Best $\beta$ (Linear): 0.4, 0.9, 0.8, 0.4 for CDs, Games, Toys, Books. Best $\beta$ (Binary): 0.2, 0.2, 0.5, 0.1 for CDs, Games, Toys, Books.
Dataset Metric Baseline (D3) IGD-T (Linear) IGD-T (Binary, ours)
CDs HR@10 0.1040 0.1072 0.1077
NDCG@10 0.0767 0.0793 0.0800
Games HR@10 0.0773 0.0800 0.0942
NDCG@10 0.0477 0.0492 0.0594
Toys HR@10 0.1029 0.1034 0.1063
NDCG@10 0.0698 0.0692 0.0719
Books HR@10 0.0315 0.0339 0.0422
NDCG@10 0.0228 0.0245 0.0312
Due to the dominance of zero-IG tokens and the non-linear nature of non-zero IG values, a binary separation with a calibrated $\beta$ provides stronger and more stable improvements than a linear mapping, across all evaluated datasets.
Appendix H Effective Hyperparameter Ranges
Our method introduces two parameters: a decoding weight $\alpha$ and a training-time weight $\beta$. Since $\alpha$ is easy to tune, we only analyze $\beta$’s sensitivity while fixing $\alpha{=}0.0$. Unless otherwise noted, the search grid for $\beta$ is $\{0.08,0.1,0.2,0.3,0.4,0.5,0.6\}$; after selecting $\beta$, $\alpha$ is tuned over $\{0.1,0.2\}$.
Optimal $\beta$. CDs: $\beta{=}0.2$ (HR@10=0.1077). Games: $\beta{=}0.2$ (HR@10=0.0942). Toys: $\beta{=}0.5$ (HR@10=0.1063). Books: $\beta{=}0.1$ (HR@10=0.0422).
Effective ranges (HR@10 and NDCG@10 >> baseline). CDs: $[0.1,0.6]$; Games: $0.1,0.4]\cup\{0.6\}$; Toys: $[0.2,0.6]$; Books: $[0.1,0.6]$.
Observation. Each dataset exhibits a broad interval where performance exceeds the baseline, making $\beta$ easy to integrate into existing methods. However, to obtain the optimal $\beta$, one still needs to search over a reasonable range.